There are three separate tunnels built side. Each is 50,000 meters long and 4ft. high. How many cubic meters of dirt had to be moved using 3.14 pi?
Answer by mananth(16949)

1 ft. = 0.3048 m
4ft = 4*0.3048=1.22m
we assume it is a cylinder of 50000m with diameter 1.22m
radius = 0.61m
volume V = pi*r^2 * h
V= 3.14 *0.61*0.61*50000
V = 58419.7 cu.m of dirt was removed
this is for 1 tunnel
you calculate for 3 tunnels.